# https://hubert0527.github.io/MALD-NeRF

### v1

@@ Abstract
Figure 1: NeRF inpainting. Given a set of posed images associated with inpainting masks, the proposed framework estimates a NeRF that renders high-quality novel views, where the inpainting region is realistic and contains high-frequency details.
@@ 1 Introduction
To address the NeRF inpainting problem, existing algorithms first leverage a 2D generative prior to inpaint the input images, then optimize a NeRF using the inpainted images. Several efforts [26, 25, 35] use the LaMa [35] model as the 2D inpainting prior. Driven by the recent success of diffusion models [10, 30, 32, 9, 4], recent work [28, 39] use the latent diffusion model [30] to further enhance the fidelity. Nevertheless, unrealistic visual appearance and incorrect geometry are still observed in the inpainted NeRFs produced by these methods.
@@
We conduct extensive quantitative and qualitative experiments on two NeRF inpainting benchmark datasets consisting of multiple challenging real-world scenes. Our proposed method, name MALD-NeRF, achieves state-of-the-art NeRF inpainting performance by marrying the merits of the Masked Adversarial learning and the Latent Diffusion model. MALD-NeRF synthesizes inpainting areas with high-frequency details and mitigates the texture shift issues. In addition, we conduct extensive ablation studies to dissect the effectiveness of each component and the effect of different loss function designs. We summarize the contributions as follows:
@@ 2 Related Work
Figure 4: 2D image inpainting comparisons. The inpainting quality of the LaMa [35] approach is less realistic compared to the LDM [30] model.
@@ 2.2 NeRF Inpainting
Recent studies on NeRF inpainting focus on inpainting individual 2D images and mitigating the 3D inconsistency problem with different solutions. NeRF-In [33] shows that simply training the NeRF with pixel reconstruction loss with inpainted images leads to blurry results. SPIn-NeRF [26] is a more recent work that proposes that replacing pixel loss with more relaxed perceptual loss enables the NeRF to reveal more high-frequency details, and the method also leverages depth predictions to supervise the NeRF geometry. [42] is a concurrent work of SPIn-NeRF, which also utilizes LaMa [35] as the image inpainter. However, as shown in Figure 4, the overall LaMa inpainting quality remains sub-optimal compared to diffusion models, leading to less compelling visual quality. InpaintNeRF360 [39] also adopts a similar strategy that relies on perceptual and depth losses. However, in our study, we found using perceptual loss does not fundamentally resolve the problem, and often leads to sub-optimal quality. Reference-guided inpainting [25] introduces a more sophisticated mechanism with careful per-view inpainting, view selection, and inpainting by referencing previous results to address the cross-view inconsistency. However, the approach is highly heuristic and sophisticated, the authors have not released either the implementation or rendering results. We have reached out to the authors to obtain visual results for comparison, but we did not receive responses. Inpaint3D [28] proposes to train a diffusion model on the RealEstate10k [48] dataset and shows high-quality scenes from a similar data distribution. However, due to the limited size of the training dataset, the method is unable to generalize to arbitrary scenes, such as the SPIn-NeRF benchmark. In contrast, we utilize a diffusion model pretrained on a large-scale internal dataset that serves as a strong prior model with outstanding generalization and shows that our method can be applied to unseen dataset distribution, such as the SPIn-NeRF benchmark.
@@ 3.2 NeRF Inpainting with Latent Diffusion Models
Masked adversarial training for inpainting regions. We do not use pixel distance losses in the inpainting region, as they are not robust to the highly diverse and 3D inconsistent inpainting results leading to blurry mist-like NeRF renderings. To address the issue, we use adversarial loss [12, 20] and the discriminator feature matching loss [41] to guide the supervision of NeRF in the inpainting regions. Specifically, we consider the patches of inpainted images as real examples, and the NeRF-rendered patches as the fake ones in the adversarial training.
However, as shown in Figure 2, the real pixels and the inpainted pixels have a textural shift, causing the discriminator to exploit such a property to recognize the real image patches. In this case, the discriminator promotes the textural discrepancy between the NeRF-rendered pixels in the reconstruction and inpainting regions. To alleviate the issue, we design a masked adversarial training scheme to hide the reconstruction/inpainting boundary on the image patches from the discriminator. For both NeRF-rendered and diffusion-inpainted images, we only keep the pixels within the inpainting mask region, and mask the pixels outside the masked region with black pixels. The design is conceptually similar to AmbientGAN [7] which trains the discriminator with corrupted inputs based on the underlying task of interest. In particular, our objective is to reduce the textural gap between the inpainting and non-inpainting renderings, instead of solving image restoration tasks. We later show that such a design indeed leads to superior inpainting geometry and performs better in quantitative evaluations in the ablation study.
@@
Total training objective. Each training iteration of MALD-NeRF consists of three steps, each step optimizes the modules with different objectives. A reconstruction step optimizing NeRF with
$L^{r}=\lambda_{\text{pix}}L_{\text{pix}}+\lambda_{\text{inter}}L_{\text{inter}}+\lambda_{\text{distort}}L_{\text{distort}}+\lambda_{\text{decay}}L_{\text{decay}}\,.$ (4)
An inpainting step optimizing NeRF with
$L^{m}=-\lambda_{\text{adv}}L_{\text{adv}}+\lambda_{fm}L_{\text{fm}}+\lambda_{\text{inter}}L_{\text{inter}}+\lambda_{\text{distort}}L_{\text{distort}}+\lambda_{\text{decay}}L_{\text{decay}}\,.$ (5)
Finally, the discriminator training step using objective
$L^{D}=L_{\text{adv}}+\lambda_{\text{GP}}L_{\text{GP}}.$ (6)
The $\lambda$’s are weighting factors of these objectives.
Iterative dataset update and noise scheduling in inpainting diffusion. In practice, directly inpainting the input images leads to inconsistency issues across various viewpoints due to the high diversity and randomness of the diffusion model. We leverage two strategies to mitigate the issue. First, we use an iterative dataset update (IDU) approach similar to [13], where we gradually update the inpainting region every $U$ iterations throughout the training. Second, the inpainting diffusion model only performs a partial DDIM [34] starting from time step $t$ based on the current NeRF rendering. The time step $t$ is determined based on the ratio of the training progress, i.e., the earlier the training is, the more noise is added to the inpainting region for denoising. Such a design aims to leverage the 3D consistency of NeRF rendering and gradually propagate such 3D consistent information to all images in the dataset. Specifically, we use the HiFA [49] scheduling that sets $t=t_{max}-(t_{max}-t_{min})*\sqrt{k/K}$, where $k$ is the current NeRF training time step, $K$ is the total iterations of the NeRF training, and $(t_{max},t_{min})=(980,20)$ is the DDPM [15] time steps used within the latent diffusion model.
@@ 4.1 Experimental Setups
Following prior work [26, 28], we resize all the images to have a long-edge size of $1008$.
@@
  • SPIn-NeRF [26]: We use the official implementation. Note that the authors do not provide the evaluation implementation. Therefore, the LPIPS scores reported in this paper differ from those presented in the SPIn-NeRF paper. Nevertheless, we have contacted the authors to ensure that the SPIn-NeRF results match the quality shown in the original paper.
  • SPIn-NeRF (LDM): We replace the LaMa [35] model to our latent diffusion model in the SPIn-NeRF approach while maintaining all default hyper-parameter settings.
@@ 5 Conclusions and Discussions
Limitations. Our framework involves an adversarial objective. Despite recent generative adversarial networks literature advancements, the framework’s performance remains highly stochastic. Also, it may not work well with low-shot NeRF reconstructions due to limited training data, or application scenarios with excessively large inpainting masks. Our framework significantly improves NeRF’s convergence by making the diffusion model generate consistent results across frames. However, our approach does not eliminate the microtextural variation caused by the stochastic denoising process. Therefore, the inpainted texture remains visually more blurry than the real-world textures.
@@ Appendix A More Visual Results
Appendix B Implementation Details

### v2

@@ Abstract
Figure 1: NeRF inpainting. Given a set of posed images associated with inpainting masks, the proposed framework estimates a NeRF that renders high-quality novel views, where the inpainting region is realistic and contains high-frequency details. Our algorithm works for both forward-facing scenes and 360 scenes, and supports both single object and multiple objects removal.
@@ 1 Introduction
To address the NeRF inpainting problem, existing algorithms first leverage a 2D generative prior to inpaint the input images, then optimize a NeRF using the inpainted images. Several efforts [28, 27] use the LaMa [37] model as the 2D inpainting prior. Driven by the recent success of diffusion models [10, 32, 34, 9, 4], recent work [30, 44] use the diffusion models to further enhance the fidelity. Nevertheless, unrealistic visual appearance and incorrect geometry are still observed in the inpainted NeRFs produced by these methods.
@@
We conduct extensive quantitative and qualitative experiments on two NeRF inpainting benchmark datasets consisting of multiple challenging real-world scenes. Our proposed method, name MALD-NeRF, achieves state-of-the-art NeRF inpainting performance by marrying the merits of the Masked Adversarial learning and the Latent Diffusion model. MALD-NeRF synthesizes inpainting areas with high-frequency details and mitigates the texture shift issues. In addition, we conduct extensive ablation studies to dissect the effectiveness of each component and loss function designs. We summarize the contributions as follows:
@@ 2.2 NeRF Inpainting
Recent studies on NeRF inpainting focus on inpainting individual 2D images and mitigating the 3D inconsistency problem with different solutions. NeRF-In [35] shows that simply training the NeRF with pixel reconstruction loss with inpainted images leads to blurry results. SPIn-NeRF [28] is a more recent work that proposes that replacing pixel loss with more relaxed perceptual loss enables the NeRF to reveal more high-frequency details, and the method also leverages depth predictions to supervise the NeRF geometry. [48] is a concurrent work of SPIn-NeRF, which also utilizes LaMa [37] as the image inpainter. InpaintNeRF360 [44] also adopts a similar strategy that relies on perceptual and depth losses. However, in our study, we found using perceptual loss does not fundamentally resolve the problem, and often leads to sub-optimal quality. Reference-guided inpainting [27] introduces a more sophisticated mechanism with careful per-view inpainting, view selection, and inpainting by referencing previous results to address the cross-view inconsistency. However, the approach is highly heuristic and sophisticated, the authors have not released either the implementation or rendering results. We have reached out to the authors to obtain visual results for comparison, but we did not receive responses. Inpaint3D [30] proposes to train a diffusion model on the RealEstate10k [54] dataset and shows high-quality scenes from a similar data distribution. However, due to the limited size of the training dataset, the method is unable to generalize to arbitrary scenes, such as the SPIn-NeRF benchmark. In contrast, we utilize a diffusion model pretrained on a large-scale internal dataset that serves as a strong prior model with outstanding generalization and shows that our method can be applied to unseen dataset distribution, such as the SPIn-NeRF benchmark. NeRFiller [47] is a concurrent work that also works updating NeRF with constrained inpainting results using diffusion prior.
@@ 3.2 NeRF Inpainting with Latent Diffusion Models
Masked adversarial training for inpainting regions. We do not use pixel distance losses in the inpainting region, as they are not robust to the highly diverse and 3D inconsistent inpainting results leading to blurry mist-like NeRF renderings. To address the issue, we use adversarial loss [13, 22] and the discriminator feature matching loss [46] to guide the NeRF in the inpainting regions. Specifically, we consider the patches of inpainted images as real examples, and the NeRF-rendered patches as the fake ones in the adversarial training.
However, as shown in Figure 2, the real pixels and the inpainted pixels have a textural shift, causing the discriminator to exploit such a property to recognize the real image patches. In this case, the discriminator promotes the textural discrepancy between the NeRF-rendered pixels in the reconstruction and inpainting regions. To alleviate the issue, we design a masked adversarial training scheme to hide the reconstruction/inpainting boundary on the image patches from the discriminator. For both NeRF-rendered and diffusion-inpainted images, we only keep the pixels within the inpainting mask region, and mask the pixels outside the masked region with black pixels. The design is conceptually similar to AmbientGAN [7] which trains the discriminator with corrupted inputs based on the underlying task of interest. In particular, our objective is to reduce the textural gap between the inpainting and non-inpainting renderings, instead of solving image restoration tasks. We later show the design indeed improves inpainting geometry and performs better in quantitative evaluations in the ablation study.
@@
Total training objective. Each training iteration of MALD-NeRF consists of three steps, each step optimizes the modules with different objectives. A reconstruction step optimizing NeRF with $L^{r}=\lambda_{\text{pix}}L_{\text{pix}}+\lambda_{\text{inter}}L_{\text{inter}}+\lambda_{\text{distort}}L_{\text{distort}}+\lambda_{\text{decay}}L_{\text{decay}}\,.$ An inpainting step optimizing NeRF with $L^{m}=-\lambda_{\text{adv}}L_{\text{adv}}+\lambda_{fm}L_{\text{fm}}+\lambda_{\text{inter}}L_{\text{inter}}+\lambda_{\text{distort}}L_{\text{distort}}+\lambda_{\text{decay}}L_{\text{decay}}\,.$ Finally, the discriminator training step using $L^{D}=L_{\text{adv}}+\lambda_{\text{GP}}L_{\text{GP}}.$ The $\lambda$’s are objective weights.
Iterative dataset update and noise scheduling in inpainting diffusion. In practice, directly inpainting the input images leads to inconsistencies across various viewpoints due to the high diversity and randomness of diffusion models. We leverage two strategies to mitigate the issue. First, we use an iterative dataset update (IDU) similar to [15], where we gradually update the inpainting region every $U$ iterations throughout the training. Second, the inpainting diffusion model only performs a partial DDIM [36] starting from time step $t$ based on the current NeRF rendering. The time step $t$ is determined based on the ratio of the training progress, i.e., the earlier the training is, the more noise is added to the inpainting region for denoising. Such a design aims to leverage the 3D consistency of NeRF rendering and gradually propagate the 3D consistent information to all images in the dataset. Specifically, we use the HiFA [55] scheduling that sets $t=t_{max}-(t_{max}-t_{min})*\sqrt{k/K}$, where $k$ is the current NeRF training time step, $K$ is the total iterations of the NeRF training, and $(t_{max},t_{min})=(980,20)$ is the DDPM [17] time steps used within the latent diffusion model.
@@ 4.1 Experimental Setups
Following [28, 30], we resize all the images to have a long-edge size of $1008$.
@@
  • FVD: To quantify the geometric consistency, we consider the image set of each scene as a video sequence, then quantify the similarity between the real and inpainted sequences. We use FVD [42] that measures both the perceptual and temporal consistency of video sequence. In order to create sufficient evaluation samples, we sub-sample each scene into multiple short sequences, each has a length of ten frames. However, the blurry results can have better temporal consistency and result in better FVD scores, therefore, the metric should be considered along with previously mentioned FID/KID.
@@
  • SPIn-NeRF [28]: Note that the authors do not provide the evaluation implementation. Therefore, the LPIPS scores reported in this paper differ from those presented in the SPIn-NeRF paper. Nevertheless, we have contacted the authors to ensure that the SPIn-NeRF results match the quality shown in the original paper.
  • SPIn-NeRF (LDM): We use our latent diffusion model in the SPIn-NeRF approach as the inpainting module while maintaining all default hyper-parameter settings.
@@ 5 Conclusions and Discussions
Limitations. Our framework involves an adversarial objective. Despite recent generative adversarial networks literature advancements, the framework’s performance remains highly stochastic. Also, it may not work well with few-shot NeRF reconstructions due to limited training data, or application scenarios with excessively large inpainting masks. Our framework significantly improves NeRF’s convergence by making the diffusion model generate consistent results across frames. However, our approach does not eliminate the microtextural variation caused by the stochastic denoising process. Therefore, the inpainted texture remains visually more blurry than the real-world textures.
@@ Appendix A More Visual Results
Appendix B Object Insertion
A few prior works [27, 47] also demonstrates applications of inpainting objects in NeRF with text prompts. However, most of these results remain preliminary without complicated object-scene interaction. In such cases, synthetic 3D objects can be inserted with depth blend afterward and later baked into NeRF with optimization. Meanwhile, directly inpainting objects with the input masks is not effective as the masks strongly constrain the object shape. In Figure 8, similar to GaussianEditor [12], we show that our framework can achieve object inpainting by first inpainting the background then followed by inserting generated objects.
Figure 8: Insert synthesized objects with depth blending.
Appendix C Training Robustness
Trial LPIPS M-LPIPS FID KID
A 0.4147 0.3056 195.47 0.0466
B 0.4138 0.3055 194.59 0.0456
C 0.4142 0.3058 191.91 0.0486
Despite we adopt adversarial loss in our optimization objectives, our algorithm is stable with R1 regularizer [24] and gradient penalty [14]. We show the robustness by reporting the performance of three individual training trials of all scenes from SPIn-NeRF dataset. To save computation with the large amount of experiments, we use 1 GPU instead of 8 GPUs per experiment.
Appendix D Further Customization
During the per-scene customization, due to repetitively using the same set of images as the finetuning data distribution, the model can sometimes memorize certain scene-dependent content. The behavior can lead to inpainting objects in the inpainting region, such as the baseball cap inpainted in Figure 5 of the main paper. Such a behavior can easily remedy by masking the unwanted objects from the scene-dependent customization. We show such a result on the right.
Appendix E Evaluate Geometric Consistency with Optical Flow-based Metric
Following [41], we evaluate the cross-frame consistency with a flow warping score. We treat the image set of each scene as an image sequence, then calculate optical flow between an image pair $I_{t}$ and $I_{t+1}$ with RAFT [40], finally, obtain $\tilde{I}_{t}$ by warping $I_{t+1}$ with the optical flow. We compute the consistency with $M_{t\rightarrow t+1}\cdot\|I_{t}-\tilde{I}_{t}\|_{1}$, where $M_{t\rightarrow t+1}$ is the disocclusion mask from time $t$ to $t+1$. We found such a metric highly depends on the flow quality predicted by RAFT, and heavily favors blurry inpainting results. In practice, we found the blurry results from all baselines outperforms the consistency scores from the real images.
Appendix F Implementation Details